Shift report 20011113 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Nov 13 2001 - 03:49:24 EST

  • Next message: Ian Bearden: "12 degrees is GOOD, was: RE: run plans and rates..."

    ------------------------------------------------------------
    Date: Tue Nov 13 03:49:24 2001, Shift: 00:00-08:00
    Supervisor: TML, with: 
    Log book#10 pages 135 - 135
    
    Shift Summary:
    ------------------------------------------------------------
    1. RHIC had RF problem when shift started. Syncronisation between AGS and RHIC. Then at 0.35 they had power supply problem, and estimated beam at 5.00. 3.35 the called and said there would be no beam before the sceduled access to the IR.
    
    2. nope
    
    3. nope 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Nov 13 2001 - 03:49:47 EST